OpenStudy (anonymous):

A circle is inscribed inside a square as shown. If graphed on the coordinate plane, the coordinates of the center of the circle would be (2, 5) and the coordinates of one other point on the circle would be (-2, 2). Find the area of the square.

OpenStudy (anonymous):

Center (2, 5): h = 2 k = 5 Point (- 2, - 2): x = - 2 y = - 2 Radius = r: r² = (x - h)² + (y - k)² r² = (- 2 - 2)² + (- 2 - 5)² r² = (- 4)² + (- 7)² r² = 16 + 49 r² = 65 r = √65 r = 8.0623 Side of Square = s: s = 2r, so s = 2(8.0623) s = 16.1245 Area of Square = A: A = s²: A = (16.1245)² A = 260 Area of Square = 260 sq. units. ¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

well there are options A) 20 B) 25 C) 40 D) 100

OpenStudy (anonymous):

ok let me recalculate

OpenStudy (anonymous):

ok sorry for not putting the options

OpenStudy (anonymous):

its ok :)

OpenStudy (anonymous):

yeah